Ramp that is 24.3 feet long reaches up to a platform from level ground. If the ramp's angle of elevation is 7.5 degrees, find the height of the platform.
Answer #1

Let P be the Height of the platform

sin(7.5)= P/24.3

P=24.3* sin(7.5)

P=24.3*0.13

P=3.159 feet